OpenStudy (danielbarriosr1):

A company builds computers. It costs $6700 to build 10 computers and $12200 to build 20 computers. Which equation models the cost C(x), as a linear function of the number of computers built, x?

OpenStudy (danielbarriosr1):

Possible answers: A. C(x)=550x-1200 B. C(x)=550x+1200 C. C(x)=1200x-550 D. C(x)=1200x+550

OpenStudy (danielbarriosr1):

@mathmale

OpenStudy (danielbarriosr1):

@acxbox22

OpenStudy (mathmale):

Daniel, this problem is much like the previous one. You are given two points on a straight line (which may not be immediately obvious); you need to find the equation of this line, doing exactly what we did last time. The two points are (10, $6700) and (20, $12200). Do you see why? Please find the slope of the line passing thru these two points.

OpenStudy (mathmale):

Once you've done that, we'll solve your equation for y, which will also be the desired function, C(x).

OpenStudy (danielbarriosr1):

m=550

OpenStudy (danielbarriosr1):

y-6700=550(x-10)

OpenStudy (danielbarriosr1):

y-6700=550x-5500

OpenStudy (mathmale):

This looks appropriate, so I'm not going to check the math (yet). Just add $6,700 to both sides of your equation. What do you get?

OpenStudy (danielbarriosr1):

y=550+1200

OpenStudy (danielbarriosr1):

That was easy

OpenStudy (mathmale):

Daniel, let's check your result. Steal one of the two points. Substitute its x into your formula. Does the resulting y agree with the y-coordinate of your point?

OpenStudy (danielbarriosr1):

6700=550(10)+1200

OpenStudy (danielbarriosr1):

6700=5500+1200

OpenStudy (mathmale):

Is that what you'd expected? If x=10, y=$6,700?

OpenStudy (danielbarriosr1):

6700=6700

OpenStudy (danielbarriosr1):

Yes

OpenStudy (mathmale):

Great. Sure looks promising! Now rewrite your equation by substituting the label C(x) for y. What do you get?

OpenStudy (danielbarriosr1):

C(x)=550x+1200

OpenStudy (danielbarriosr1):

Got it... thank you

OpenStudy (mathmale):

Got it! That's your desired formula! Cool! I need to get off the 'Net now, but very much look forward to working with you again soon! All the best to you, as always.

OpenStudy (danielbarriosr1):

I also need to leave... Same to you